OpenStudy (anonymous):

Evaluate the integral 6e^(3.4x) with limits (30,50)

OpenStudy (lgbasallote):

hmm looks familiar...wasnt able to solve it?

OpenStudy (anonymous):

nope :(

OpenStudy (lgbasallote):

why not?

OpenStudy (anonymous):

not sure guess i couldnt figure it out. either that or i am incorrectly entering answers like 2.3745 x 10^76 into my online homework

OpenStudy (anonymous):

if i even had the right answer

OpenStudy (lgbasallote):

it was \[\huge 1.77e^{17x/6}\] right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now substitute 50 and 30

OpenStudy (lgbasallote):

\[\huge 1.77 e^{\frac{17 \times 50}{5}} - 1.77e^{\frac{17 \times 30}{3}}\]

sam (.sam.):

Yes, \[\int\limits_{30}^{50}6e^{3.4x}=[\frac{6e^{3.4x}}{3.4}]_{30}^{50}\]

OpenStudy (lgbasallote):

derp...that should be over 5 not 3

sam (.sam.):

Power rule will do here

OpenStudy (lgbasallote):

\[\huge 1.77 e^{17 \times 10} - 1.77 e^{17 \times 6}\] i just canceled the thingies...did you get this?

OpenStudy (lgbasallote):

lol big number

OpenStudy (anonymous):

im still getting 2.386515463 x 10^75 soo is that right?

OpenStudy (lgbasallote):

that's not what wolfram got

OpenStudy (anonymous):

what did wolfram get

OpenStudy (lgbasallote):

you may have had a syntax error http://www.wolframalpha.com/input/?i=1.77+x+e%5E170+-+1.77+x+e%5E102

OpenStudy (anonymous):

oh yeah well my hw is saying that that answer isnt correct either grrr

OpenStudy (lgbasallote):

WHAT?!!!

sam (.sam.):

should be http://www.wolframalpha.com/input/?i=1.77+%28+e^170%29+-%28+1.77%29+%28+e^102%29 I think you put the 'times' incorrectly

OpenStudy (lgbasallote):

im right o.O

OpenStudy (lgbasallote):

well a little inaccurate but still!!

OpenStudy (anonymous):

wait so witch should i put in as my answer?

sam (.sam.):

Both are the same http://www.wolframalpha.com/input/?i=int+6e^%283.4x%29+from+30+to+50 http://www.wolframalpha.com/input/?i=1.77+%28+e^170%29+-%28+1.77%29+%28+e^102%29 \[\int\limits_{30}^{50} 6 e^{3.4 x} \, dx=1.19326\times 10^{74}\]

OpenStudy (anonymous):

do you know any other alternative forms for that answer?